# apndxAdditional References A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability

### v1

@@ 4.1.3. Protocol Integration and Compatibility with Existing Blockchains
Although based on blockchain principles, most second-generation privacy-preserving cryptocurrencies are not fully compatible with baseline systems. We analyze how they interact with existing digital currencies, whether through compatibility, non-compatible extensions, or entirely different approaches.
@@
Mixing protocols propose a supplementary services for existing blockchains, in particular for Bitcoin, and are specifically designed to be fully compatible (e.g., (Ruffing et al., 2016; Valenta and Rowan, 2015; Qin et al., 2023; Ruffing and Moreno-Sanchez, 2017)). Another category of compatible solutions leverages blockchain programmability to introduce privacy-enhanced currencies through overlays (e.g., (Bünz et al., 2020; Baudet et al., 2023)). For instance, Zether (Bünz et al., 2020) employs a smart contract on Ethereum to store and manage private accounts, whereas Zef (Baudet et al., 2023) introduces opaque coins in the form of off-chain certificates exchanged between Ethereum user accounts.
@@
Most privacy-enhancing currency proposals result in non-compatible extensions of existing blockchain protocols. For example, Zerocoin (Miers et al., 2013) is an overlay over Bitcoin that allows user to convert bitcoins in zerocoins. A zerocoin can be proved (using ZK techniques) to originate from a valid and unspent bitcoin, without revealing the link to the original bitcoin. Zerocoin does not provide full anonymity: it obscures coin traceability, but it does not hide the number of transactions, recipient addresses, and balances. Building on Zerocoin, Zerocash (Ben Sasson et al., 2014) enhances privacy by concealing both sender and recipient addresses, supporting arbitrary transaction values, and enabling change handling. Both Zerocoin and Zerocash require new transaction types and payment semantics, and break compatibility with Bitcoin. Garman et al. (Garman et al., 2017) extends Zerocash with no backward compatibility, by introducing policy enforcement, selective user tracing, and tracing of tainted coins—features that further diverge from Bitcoin’s existing design. Yu et al. (Yu et al., 2024) propose an overlay on Bitcoin for anonymous yet regulated transactions using a coin backed by bitcoins. Their approach relies on three centralized entities that cooperate to enable batch linkability and payer tracing while preserving privacy. These changes are not backward-compatible with Bitcoin.
@@
Most approaches present greenfield solutions for privacy-native cryptocurrencies (e.g., (Van Saberhagen, 2013; Poelstra, 2016; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020)), often extending open-source solutions (e.g., (Li et al., 2021; Papadoulis et al., 2024)). Both centralized (e.g., (Cecchetti et al., 2017; Zhang et al., 2019)) and decentralized architectures (e.g., (Narula et al., 2018; Androulaki et al., 2020; Lin et al., 2020; Ma et al., 2021; Goodell et al., 2023; Elfadul et al., 2024)) for privacy-native solutions have been proposed, with a sensible preference for the latter. Early approaches like CryptoNote (Van Saberhagen, 2013), MimbleWimble (Poelstra, 2016), and Monero (Noether et al., 2016) were among the first to address Bitcoin’s privacy limitations by still proposing fully decentralized alternatives. CryptoNote (Van Saberhagen, 2013) enhances sender and receiver anonymity through one-time ring signatures and NIZK proofs, inspiring cryptocurrencies like ByteCoin, DigitalNote, and Aeon (Amarasinghe et al., 2019). MimbleWimble (Poelstra, 2016) improves privacy and scalability by aggregating confidential transactions instead of storing individual transaction data. Cryptocurrencies building on MimbleWimble include Grin444https://grin.mw/ and Beam555https://www.beam.mw/. Monero (Noether et al., 2016) strengthens privacy with stealth addresses, ring signatures, and confidential transactions, while also integrating network anonymity protocols.
Other approaches propose partially decentralized solutions (e.g., (Wüst et al., 2019)), combine concepts from centralized mixers (e.g., (Fauzi et al., 2019)), or consider hierarchical bank-intermediated settings (e.g., (Cecchetti et al., 2017; Narula et al., 2018)). Wüst et al. (Wüst et al., 2019) propose PRCash to strike a trade-off between performance (fast payments), user privacy, and regulatory oversight. Basically, PRCash includes two centralized entities, which respectively issue money and oversee compliance, and a permissioned blockchain, which records MimbleWimble-like transactions. Quisquis (Fauzi et al., 2019) builds on the concept of mixing to redistribute wealth: each transaction updates a bunch of other accounts on chain, thus preventing addresses from appearing multiple times on the blockchain, reducing the risk of de-anonymization and linkability. Quisquis ensures integrity through ZKPs while maintaining privacy using commitments and updatable public keys. Cecchetti et al. (Cecchetti et al., 2017) propose Solidus, a hierarchical, bank-intermediated systems where a small number of banks manage transactions of on-chain assets on behalf of a large number of users. This setup ensures transaction-graph confidentiality: sender and receiver of a transaction cannot be publicly identified666While banks can identify their respective clients, other entities only learn the identities of the banks., even by pseudonyms.
@@ 4.1.4. Trust Assumptions
We now focus on the degree of trust required, such as reliance on a trusted third party, to facilitate secure transactions. Most solutions for digital and cryptocurrencies do not usually rely on trusted parties for confidentiality (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Li et al., 2021; Ma et al., 2021; Guan et al., 2022; Tomescu et al., 2022; Baudet et al., 2023; Nosouhi et al., 2023; Sarencheh et al., 2025; Wang et al., 2023)), even when they propose a centralized system (e.g., (Chaum, 1983; Chaum et al., 2021; Wüst et al., 2022)).
The solutions that introduce trust assumptions do it for different purposes. The approaches that use zk-SNARK as ZKP require a trusted setup to initialize and correctly share some cryptographic parameters (e.g., (Miers et al., 2013; Garman et al., 2017; Ben Sasson et al., 2014; Xue et al., 2023)). A few works require a trusted central bank to issue coins that can be later exchanged anonymously (e.g., (Goodell et al., 2023; Friolo et al., 2024)). Elfadul et al. (Elfadul et al., 2024) assume the presence of a regulatory supervision that grants permissions to users to perform transaction; this centralized authority could asymmetrically revoke such authorization. Contributions that explore special forms of auditability related to revocable anonymity (e.g. (Wüst et al., 2019; Lin et al., 2020; Xue et al., 2023; Yu et al., 2024)), trapdoor (e.g., (Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Elfadul et al., 2024)), or user tracing (e.g., (Garman et al., 2017)) require trust in the tracing authority that uncovers transaction details only when necessary. PEReDi (Kiayias et al., 2022) reduces trust requirements by relying on quorum of centralized entities that must agree to perform specific operations, such as decrypting transaction information.
@@ 4.1.5. Interaction Models
Privacy-preserving digital currencies differ in the extent to which transactions require direct or indirect interaction between the payer and the payee. Interactive payment protocols require active cooperation between payer and payee during the transaction process. This is typical in many account-based models where the recipient must be involved in constructing or approving the updated account state, often through ZKPs or digital signatures. For instance, PEReDi (Kiayias et al., 2022) and Platypus (Wüst et al., 2022) enforce synchronized interaction: both parties participate in updating the ledger with proofs and commitments, ensuring mutual accountability and preventing attacks such as dusting or replay. Similarly, Chaumian token-based systems often require the payee to claim tokens immediately after receipt to prevent double spending (e.g., (Dold, 2019)). In contrast, non-interactive payment protocols allow the payer to unilaterally construct a transaction and broadcast it to the network. Many token-based systems, especially those based on unspent tokens, adopt this strategy (e.g., (Ben Sasson et al., 2014; Van Saberhagen, 2013; Noether et al., 2016; Poelstra, 2016)). This model improves user convenience and enables asynchronous payments, but may require periodic ledger scanning or local computation by the receiver777Off-chain communication can alleviate the burden of scanning the ledger, but current protocols lack this feature.. Some account-based protocols implement delayed claiming to approximate this non-interactive behavior. Sarencheh et al. (Sarencheh et al., 2025) implement a non-interactive variant of PEReDi, which works in two stages. In the transfer stage, the sender interacts with a set of issuers to update his account and provide information in ZK. The claim stage credits the receiver account, exploiting information enclosed in ZKP by the sender. While interactive designs enable rich policy enforcement and conditional privacy, non-interactive models better suit high-latency or asynchronous networks.
@@ 4.2. Value Representation
The value representation dimension examines the approaches used to represent the currency, the user balance, and the exchange of value. For the purpose of this survey, we investigate these aspects with a specific focus on their relation with privacy-related features.
@@ 4.2.1. Ledger Data Models
Digital currencies and cryptocurrencies follow three main approaches for representing and tracking ownership of digital assets: accounts, tokens, and tabular ledger.
@@
Account-based ledgers associate a balance with each account address, which is debited or credited by transactions. This simplifies state management and enables simple definitions of policies based on cumulative balances. Nonetheless, it requires safeguards against replay and double-spending attacks, often implemented by storing a sequence number (or nonce) in each account and updating it with every transaction. For example, Ethereum (Buterin et al., 2013) associates to each account888Ethereum also defines standards to represent fungible tokens (ERC-20) and non-fungible tokens (ERC-721). ERC-20 can be used to design privacy-preserving currencies (e.g., (Bünz et al., 2020)). a nonce (indicating the number of sent transactions), a balance, and two hashes referencing the account’s storage contents and associated code. Since the balance of an account is explicitly represented, obtaining anonymity and unlinkability is challenging, especially in systems where users cannot freely generate multiple addresses (e.g., (Guan et al., 2022)). Privacy-preserving account-based solutions replace plain-text balances with commitments, achieving confidentiality of balances and transactions (e.g., (Fauzi et al., 2019; Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022; Sarencheh et al., 2025)). Liu et al. (Liu et al., 2022) propose an account-based CBDC with sender and receiver anonymity but with no unlinkability. The system follows a two-tier model where the central bank can identify users via encrypted identifiers, while commercial banks cannot. To guarantee anonymity while ensuring unlinkability, most of the approaches change users’ account address at each transaction and use an authenticator to prove correctness. The authenticator can be implemented, e.g., using digital signatures by trusted or centralized authorities (e.g., (Kiayias et al., 2022)) or ZKPs (e.g., (Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)). Platypus (Wüst et al., 2022) rely on ZKPs; conversely, in PEReDi (Kiayias et al., 2022), account states must be (blindly) signed by a quorum of maintainers, and since each transaction reveals prior account snapshots to construct new authenticators, randomized signatures are used to prevent linkability. Quisquis (Fauzi et al., 2019) and AQQUA (Papadoulis et al., 2024) use updatable public keys to update accounts in a decentralized manner. The account-based model lends itself well to auditability, as it allows simple definition of policies on holding limits (e.g., (Gross et al., 2021; Wüst et al., 2022; Papadoulis et al., 2024)) or spending and receiving limits (e.g., (Papadoulis et al., 2024)). For example, the account state by Gross et al. (Gross et al., 2021) includes an accumulator of all spending in the current epoch, enabling policies to rely on this information. AQQUA (Papadoulis et al., 2024), instead, includes commitments to three different counters (i.e., balance, total amounts sent and received). Conversely, Platypus (Wüst et al., 2022) leaves the specification of the additional information to be application dependent.
@@
Token-based ledgers do not maintain explicit balances per address, but instead track individual tokens. The user’s balance can be derived from the sum of all owned tokens. This model enables parallel validation of user transactions and stronger privacy guarantees (as balances are not explicitly materialized) but introduces management complexity, particularly when users distribute their coins across multiple addresses for added privacy. There are two main types of token structures: Chaumian tokens and unspent tokens. Although both represent digital value, they differ in structure, spending mechanisms, privacy guarantees, and centralization.
@@
Chaumian tokens (Chaum, 1983) use blind signatures and were first deployed in centralized, bank-issued electronic cash schemes. When issuing a token, the bank blindly signs a random value provided by the payer, allowing the payer to later unblind it for spending. Blind signatures prevent the bank from linking withdrawals to subsequent transactions, thus ensuring sender anonymity. These tokens rely on a trusted authority to issue and verify their validity; to this end, the authority must hold an ever increasing list of spent tokens that might introduce scalability challenges. Some implementations introduce expiration to limit circulation (e.g., (Dold, 2019; Lovejoy et al., 2023)). The Chaumian model inspired decentralized privacy-preserving cryptocurrencies (e.g., Monero (Noether et al., 2016), Zerocoin (Miers et al., 2013), and Zerocash (Ben Sasson et al., 2014)), which eliminate the need for a central issuer by leveraging a combination of different cryptographic techniques. For example, Zerocash supports transparent (like Bitcoin) and shielded transactions, with the latter hiding sender, recipient, and amount. Shielded payments encrypt these values with one-time keys and use zk-SNARK proofs to validate inputs-outputs equality, concealing all on-chain details.
@@
Unspent tokens, introduced in Bitcoin (Nakamoto, 2008), represent transaction outputs that can be used as inputs for future transactions, which in turn create new unspent tokens. As such, they are inherently traceable unless PETs are used (Baldimtsi et al., 2024). Unspent tokens are issued in a decentralized manner, usually as result of the mining process. Their core attributes include an identifier, value (or its commitment), and ownership information. Their recording in a public, transparent ledger prevents double-spending but complicates privacy. Privacy-focused cryptocurrencies employ stealth addresses, ring signatures, confidential transactions, and ZKPs to obscure transaction details and unlink payments (e.g., (Noether et al., 2016; Poelstra, 2016; Miers et al., 2013; Ben Sasson et al., 2014))—these properties are discussed in more detail later. Tokens can also include auxiliary data to support the evaluation of specific policies (e.g., (Xue et al., 2023; Yu et al., 2024)). For example, Xue et al. (Xue et al., 2023) include information to prove in ZK that the total amount of the cryptocurrency transferred and the transaction frequency in a time period is limited. For a detailed discussion of auditability in token-based systems, see Section 4.4.
@@
Narula et al. (Narula et al., 2018) propose zkLedger, a third approach tailored to a specific use case: a bank-intermediated CBDC. It uses a tabular ledger where rows represent transactions and columns represent banks. Each cell contains a Pedersen commitment to hide values, while enabling cryptographic verification. Transactions include ZKPs to prove balance, asset ownership, and consistency with audit records. This structure supports rich privacy-preserving auditing: the system includes an auditor who can issue a rich set of auditing queries—using primitives like sums, moving averages, variance, standard deviation, ratios—and receive answers that are provably consistent with the ledger. As downside, this structure does not scale well with many banks and transactions, making it suitable for small, privacy-focused networks.

### v2

@@ 4.1.3. Protocol Integration and Compatibility with Existing Blockchains
Second-generation privacy-preserving payment systems vary in their integration with existing (blockchain) infrastructures: some remain compatible, others introduce non-compatible extensions, and some adopt entirely new designs.
@@
Mixing protocols propose a supplementary services for existing blockchains (mostly for Bitcoin), and are specifically designed to be fully compatible (e.g., (Ruffing et al., 2016; Valenta and Rowan, 2015; Qin et al., 2023; Ruffing and Moreno-Sanchez, 2017)). Another category of compatible solutions leverages blockchain programmability to introduce privacy-preserving overlays (e.g., (Bünz et al., 2020; Baudet et al., 2023)). For instance, Zether (Bünz et al., 2020) uses a smart contract on Ethereum to store and manage private accounts; Zef (Baudet et al., 2023) introduces opaque coins in the form of off-chain certificates exchanged between Ethereum user accounts.
@@
Most privacy-enhancing proposals are non-compatible with the blockchain protocol they extend. For example, Zerocoin (Miers et al., 2013) is an overlay over Bitcoin that allows user to convert bitcoins in zerocoins. A zerocoin can be proved (in ZK) to originate from a valid and unspent bitcoin, without explicitly revealing such link. Zerocoin does not provide full anonymity: it obscures coin traceability, but it does not hide the number of transactions, recipient addresses, and balances. Building on Zerocoin, Zerocash (Ben Sasson et al., 2014) enhances privacy by concealing both sender and recipient addresses, supporting arbitrary transaction values, and enabling change handling. Both Zerocoin and Zerocash require new transaction types and payment semantics, and break compatibility with Bitcoin. Garman et al. (Garman et al., 2017) extends Zerocash with no backward compatibility, by introducing policy enforcement and tracing of users and tainted coins—features that further diverge from Bitcoin’s existing design. Yu et al. (Yu et al., 2024) propose an overlay on Bitcoin for anonymous yet regulated transactions using a coin backed by bitcoins. Their approach relies on three centralized entities that cooperate to enable batch linkability and payer tracing while preserving privacy. These changes are not backward-compatible with Bitcoin.
@@
Most approaches present greenfield solutions for privacy-preserving payments or asset exchanges (e.g., (Van Saberhagen, 2013; Poelstra, 2016; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020)), often extending open-source solutions (e.g., (Li et al., 2021; Papadoulis et al., 2024)). Both centralized (e.g., (Cecchetti et al., 2017; Zhang et al., 2019)) and decentralized architectures (e.g., (Narula et al., 2018; Androulaki et al., 2020; Lin et al., 2020; Ma et al., 2021; Goodell et al., 2023; Elfadul et al., 2024)) exist, with a sensible preference for the latter. Early approaches like CryptoNote (Van Saberhagen, 2013), MimbleWimble (Poelstra, 2016), and Monero (Noether et al., 2016) address the Bitcoin’s privacy limitations by still proposing fully decentralized alternatives. CryptoNote (Van Saberhagen, 2013) enhances sender and receiver anonymity through one-time ring signatures and NIZK proofs, inspiring solutions like ByteCoin, DigitalNote, and Aeon (Amarasinghe et al., 2019). MimbleWimble (Poelstra, 2016) improves privacy and scalability by aggregating confidential transactions. Grin555https://grin.mw/ and Beam666https://www.beam.mw/ build on MimbleWimble ideas. Monero (Noether et al., 2016) strengthens privacy with sender, receiver, and network anonymity as well as confidential transactions.
Other approaches propose partially decentralized solutions (e.g., (Wüst et al., 2019)), combine concepts from centralized mixers (e.g., (Fauzi et al., 2019)), or consider hierarchical bank-intermediated settings (e.g., (Cecchetti et al., 2017; Narula et al., 2018)). Wüst et al. (Wüst et al., 2019) propose PRCash to strike a trade-off between fast payments, user privacy, and regulatory oversight: it includes two centralized entities for issuance and oversight and a permissioned blockchain, with MimbleWimble-like transactions. Quisquis (Fauzi et al., 2019) builds on the concept of mixing to redistribute wealth: each transaction updates a bunch of other accounts on chain, thus preventing addresses from appearing multiple times on the ledger, reducing the risk of de-anonymization and linkability. Cecchetti et al. (Cecchetti et al., 2017) propose Solidus, a hierarchical, bank-intermediated systems where a small number of banks manage transactions of on-chain assets on behalf of a large number of users; this setup ensures transaction-graph confidentiality.777While banks can identify their respective clients, other entities only learn the identities of the banks.
@@ 4.1.4. Trust Assumptions
We now focus on the degree of trust required to facilitate secure transactions. Most solutions for digital payments do not rely on trusted parties for confidentiality (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Li et al., 2021; Ma et al., 2021; Guan et al., 2022; Tomescu et al., 2022; Baudet et al., 2023; Nosouhi et al., 2023; Sarencheh et al., 2025; Wang et al., 2023)), even when they propose a centralized system (e.g., (Chaum, 1983; Chaum et al., 2021; Wüst et al., 2022)).
The solutions that introduce trust assumptions do it for different purposes. The approaches using zk-SNARK as ZKP require a trusted setup to initialize and correctly share some cryptographic parameters (e.g., (Miers et al., 2013; Garman et al., 2017; Ben Sasson et al., 2014; Xue et al., 2023)). A few works require a trusted central bank to issue coins that can be later exchanged anonymously (e.g., (Goodell et al., 2023; Friolo et al., 2024)). Elfadul et al. (Elfadul et al., 2024) assume the presence of a regulatory supervision that grants permissions to users to perform transaction; this centralized authority could asymmetrically revoke such authorization. Contributions that explore special forms of auditability related to revocable anonymity (e.g. (Wüst et al., 2019; Lin et al., 2020; Xue et al., 2023; Yu et al., 2024)), trapdoor (e.g., (Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Elfadul et al., 2024)), or user tracing (e.g., (Garman et al., 2017)) require trust in the tracing authority that uncovers transaction details only when necessary. PEReDi (Kiayias et al., 2022) reduces trust requirements by relying on quorum of centralized entities that must agree to perform specific operations (such as decrypting transactions).
@@ 4.1.5. Interaction Models
Privacy-preserving digital payment systems differ in whether transactions require direct interaction between payer and payee. Interactive payment protocols need active cooperation during the transaction process. This is common in account-based models where recipients help in constructing or approving the updated account state, via ZKPs or digital signatures. For example, PEReDi (Kiayias et al., 2022) and Platypus (Wüst et al., 2022) enforce synchronized interaction: both parties participate in updating the ledger with proofs and commitments, ensuring mutual accountability and preventing attacks like dusting or replay. Similarly, Chaumian token systems often require immediate token claiming to prevent double spending (e.g., (Dold, 2019)). In contrast, non-interactive payment protocols let the payer unilaterally create and broadcast transactions to the network. Many token-based systems, especially unspent token models, follow this approach (e.g., (Ben Sasson et al., 2014; Van Saberhagen, 2013; Noether et al., 2016; Poelstra, 2016)). This improves user convenience and enables asynchronous payments, although receivers may need periodic ledger scanning or local computation.888Off-chain communication can alleviate the burden of scanning the ledger, but current protocols lack this feature. Some account-based protocols approximate non-interactive behavior via delayed claiming. For instance, Sarencheh et al. (Sarencheh et al., 2025) propose a two stage and non-interactive variant of PEReDi: first, the sender interacts with issuers to update his account (transfer stage); then, the receiver claims funds using enclosed ZKP and credits his account (claim stage). Overall, interactive designs enable rich policy enforcement and conditional privacy, whereas non-interactive models better suit high-latency or asynchronous networks.
@@ 4.2. Value Representation
The value representation dimension examines how balances and transfers are modeled, focusing in particular on their privacy implications.
@@ 4.2.1. Ledger Data Models
Digital payment systems and decentralized ledgers follow three main approaches for representing and tracking ownership of digital assets: accounts, tokens, and tabular ledger.
@@
Account-based ledgers associate a balance with each account address, which is debited or credited by transactions. This simplifies state management and enables simple definitions of policies based on cumulative balances. Nonetheless, it requires safeguards against replay and double-spending attacks, often implemented by storing a sequence number (or nonce) in each account and updating it with every transaction. Since the balance of an account is explicitly represented, obtaining anonymity and unlinkability is challenging, especially in systems where users cannot freely generate multiple addresses (e.g., (Guan et al., 2022)). Privacy-preserving account-based solutions replace plain-text balances with commitments, achieving confidentiality of balances and transactions (e.g., (Fauzi et al., 2019; Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022; Sarencheh et al., 2025)). Liu et al. (Liu et al., 2022) propose an account-based CBDC with sender and receiver anonymity but with no unlinkability. The system follows a two-tier model where the central bank can identify users via encrypted identifiers, while commercial banks cannot. To guarantee anonymity while ensuring unlinkability, most of the approaches change users’ account address at each transaction and use an authenticator to prove correctness. The authenticator can be implemented, e.g., using digital signatures by trusted or centralized authorities (e.g., (Kiayias et al., 2022)) or ZKPs (e.g., (Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)). Platypus (Wüst et al., 2022) rely on ZKPs; conversely, in PEReDi (Kiayias et al., 2022), account states must be (blindly) signed by a quorum of maintainers, and since each transaction reveals prior account snapshots to construct new authenticators, randomized signatures are used to prevent linkability. Quisquis (Fauzi et al., 2019) and AQQUA (Papadoulis et al., 2024) use updatable public keys to update accounts in a decentralized manner. The account-based model lends itself well to auditability, as it allows simple definition of policies on holding limits (e.g., (Gross et al., 2021; Wüst et al., 2022; Papadoulis et al., 2024)) or spending and receiving limits (e.g., (Papadoulis et al., 2024)). Gross et al. (Gross et al., 2021) includes, in the account state, an accumulator of all spending in the current epoch, enabling policies on this information. AQQUA (Papadoulis et al., 2024) includes commitments to three different counters (i.e., balance, total amounts sent and received). Conversely, Platypus (Wüst et al., 2022) leaves the specification of the additional information to be application dependent.
@@
Token-based ledgers do not maintain explicit balances per address, but instead track individual tokens. The user’s balance can be derived from the sum of all owned tokens. This model enables parallel validation of user transactions and stronger privacy guarantees (as balances are not explicitly materialized) but introduces management complexity, particularly when users distribute their coins across multiple addresses for added privacy. There are two main types of token structures: Chaumian tokens and unspent tokens.
@@
Chaumian tokens (Chaum, 1983) use blind signatures and were first deployed in centralized, bank-issued electronic cash schemes. When issuing a token, the bank blindly signs a random value provided by the payer, allowing the payer to later unblind it for spending (ensuring sender anonymity). These tokens rely on a trusted authority to issue and verify their validity; to this end, the authority must hold an ever increasing list of spent tokens that might introduce scalability challenges. Some implementations introduce expiration to limit circulation (e.g., (Dold, 2019; Lovejoy et al., 2023)). Chaumian-inspired decentralized privacy-preserving systems (e.g., Monero (Noether et al., 2016), Zerocoin (Miers et al., 2013), and Zerocash (Ben Sasson et al., 2014)) eliminate the need of a central issuer using a combination of different cryptographic techniques. For example, Zerocash supports transparent (like Bitcoin) and shielded transactions, with the latter hiding sender, recipient, and amount. Shielded payments encrypt these values with one-time keys and use zk-SNARK proofs to validate inputs-outputs equality, concealing all on-chain details.
@@
Unspent tokens, introduced in Bitcoin (Nakamoto, 2008), represent transaction outputs that can be used as inputs for future transactions, which in turn create new unspent tokens. As such, they are inherently traceable unless PETs are used (Baldimtsi et al., 2024). Unspent tokens are issued in a decentralized manner, usually as result of the mining process. Their core attributes include an identifier, value (or its commitment), and ownership information. Their recording in a public, transparent ledger prevents double-spending but complicates privacy. Privacy-focused crypto-asset systems employ stealth addresses, ring signatures, confidential transactions, and ZKPs to obscure transaction details and unlink payments (e.g., (Noether et al., 2016; Poelstra, 2016; Miers et al., 2013; Ben Sasson et al., 2014))—these properties are discussed in more detail later. Tokens can also include auxiliary data to support the evaluation of specific policies (e.g., (Xue et al., 2023; Yu et al., 2024)). For example, Xue et al. (Xue et al., 2023) include information to prove in ZK that the total amount of value transferred and the transaction frequency in a time period is limited. For a detailed discussion of auditability in token-based systems, see Section 4.4.
@@
Narula et al. (Narula et al., 2018) propose zkLedger, a third approach tailored to a specific use case: a bank-intermediated CBDC. It uses a tabular ledger where rows represent transactions and columns represent banks. Each cell contains a Pedersen commitment to hide values, while enabling cryptographic verification. Transactions include ZKPs to prove balance, asset ownership, and consistency with audit records. This structure supports rich privacy-preserving auditing: the system includes an auditor who can issue a rich set of auditing queries—using primitives like sums, moving averages, variance, standard deviation, ratios—and receive answers that are provably consistent with the ledger. As downside, this structure does not scale well with many banks and transactions.